Software to detect fraud in consumer phone cards tracks the number of metropolitan areas where calls originate each day. It is found that 3% of the legitimate users originate calls from two or more metropolitan areas in a single day. However, 30% of fraudulent users originate calls from two or more metropolitan areas in a single day. The proportion of fraudulent users is 0.0198%. If the same user originates calls from two or more metropolitan areas in a single day, what is the probability that the user is fraudulent? Round your answer to six decimal places (e.g. 98.765432).
Answer #1

Let assume F=fraudulent users

And M= users originate calls from two or more metropolitan areas in a single day.

Given data are p (F)= 0.0198%= 0.000198

Ans p (F')=1-0.000198= 0.999802

P(M/F)=30%=0.30

P (M/F')=3%=0.03

Let find p (F/M)=If the same user originates calls from two or more metropolitan areas in a single day, then the probability that the user is fraudulen

According to Bay's therom we know that

P (F/M)= P(F) P(M/F) / [ P(F) P(M/F) + P(F') P(M/F')]

   = (0.000198 × 0.30) / [ (0.000198 × 0.30) + (0.999802 ×0.03)]

=0.0000594 / ( 0.0000594 + 0.02999406)

= 0.0000594 / 0.03005346

=0.001976477

=0.001976